OpenStudy (anonymous):

Write the equation in slope-intercept form. What are the are the slope and y-intercept?-3x-10y=7

OpenStudy (anonymous):

a.y=-3/10x-7/10;slope:-3/10;y-intercept:-7/10 b.y=3/10x-7/10;slope:3/10;y-intercept:-7/10 c.y=-3/10x+7/10;slope:-3/10;y-intercept:7/10 d.y=3/10x+7/10;slope:3/10;y-intercept:7/10

OpenStudy (anonymous):

can some one please explain or tell how to do this.